Solve the equation by completing the square: 3x2 - 24x = 27.
Enter the largest value of the coordinates. Use a number only.
This is a question from my math homework. I made it to 3(x^2-24x)-27, but now I'm stuck... Please help!

Answer by Alan3354(69443)

Solve the equation by completing the square: 3x2 - 24x = 27.
Enter the largest value of the coordinates. Use a number only.
This is a question from my math homework. I made it to 3(x^2-24x)-27, but now I'm stuck... Please help!
-------------
I made it to 3(x^2-24x)-27
That's no longer an equation, there's no equal sign.
3(x^2-24x)-27 = 0 is an equation.
x^2-24x-27 = 0
---------------
Can you finish it?